### v1

@@ 2 The Positive Semidefinite (PSD) Cone
The group $\mathrm{GL}(n,\mathbb{Z})$ embeds into $\mathrm{GL}(N,\mathbb{Z})$ as follows. Given a matrix $U\in\mathrm{GL}(n,\mathbb{Z})$ and any $X\in\mathcal{S}^{n}(\mathbb{Z})$, we define the action $U\cdot X\mathrel{\mathop{\ordinarycolon}}=UXU^{\mathsf{T}}$. This action is a linear map and takes integer points in $\mathbb{Z}^{N}$ to integer points, and thus can be represented by the multiplication with a matrix in $\mathrm{GL}(N,\mathbb{Z})$. It is well-known that this group $\mathrm{GL}(n,\mathbb{Z})$ is finitely generated [23]. For the convenience of discussion, we still use the matrix $U\in\mathrm{GL}(n,\mathbb{Z})$ to denote this matrix multiplication in the subgroup of $\mathrm{GL}(N,\mathbb{Z})$.
@@ 2.1 Lemmas for $n\leq 5$ and $n=6$
For any matrix $X\in\mathcal{S}^{n}(\mathbb{R})$, we can define a convex set $C(X)\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0\}$. Since $X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0$ if and only if, for any $\mathbf{v}\in\mathbb{R}^{n}$, $|\mathbf{v}^{\mathsf{T}}\mathbf{x}|^{2}\leq\mathbf{v}^{\mathsf{T}}X\mathbf{v}$, we see that $C(X)$ is a compact convex set that is symmetric about the origin. This provides another equivalent formulation of the integer rank-1 decomposition.
Proposition 2
For $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, $X$ is sporadic if and only if $C(X)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
This provides a geometric perspective to our problem. Note that the set $C(X)$ is a (possibly degenerate) ellipsoid because
@@
by the positive semidefiniteness of Schur complements, where $X^{\dagger}$ denotes the pseudoinverse of $X$. In the case where $\det(X)>0$,
$C(X)=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}\text{ with }\mathrm{vol}(C(X))=V_{n}\sqrt{\det(X)}$
where $\mathrm{adj}(X)$ is the adjugate of $X$ satisfying $\mathrm{adj}(X)=\det(X)X^{-1}$ and $V_{n}\mathrel{\mathop{\ordinarycolon}}=\pi^{n/2}/\Gamma(\frac{n}{2}+1)$ is the volume of the unit $n$-ball. Note that $C(X)$ is not necessarily full-dimensional if $X$ is rank deficient. The degenerate case for rank 1 is characterized by the following proposition.
Proposition 3
@@
From Proposition 3, we can directly prove the case for $n=2$ using Minkowski’s first Theorem (for example, see [11]): if $\mathrm{vol}(C(X))>4$, then there is a nonzero integral point.
Proposition 4
@@
In the general degenerate case, we can reduce the problem to one involving full-rank matrices of some lower dimension.
Lemma 2
@@
$\gamma_{n}\mathrel{\mathop{\ordinarycolon}}=\left(\max_{A\succ 0}\frac{\lambda_{1}(A)}{(\det(A))^{\frac{1}{n}}}\right)^{n},\text{ where }\lambda_{1}(A)=\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}A\mathbf{x}).$
Remark 1
@@
$n$
2
3
4
5
6
7
8
24
$\gamma_{n}$
$\frac{4}{3}$
2
2
8
$\frac{64}{3}$
64
256
$4^{24}$
Proof (for Lemma 1)
The case $n=1$ follows from Proposition 3. We will show that $C(X)\cap\mathbb{Z}^{n}\neq\{\mathbf{0}\}$ for $2\leq n\leq 5$, where $C(X)=\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}$. By the definition of the Hermite constant, we have
@@
For $n=2,3,4,5$, we have $\frac{n^{n}}{(n-1)^{n-1}}>\gamma_{n}$. By taking the derivative with respect to $\det(X)$, we know that $\frac{(\det(X)+1)^{n}}{(\det(X))^{n-1}}\geq\frac{n^{n}}{(n-1)^{n-1}}$. Thus, $\gamma_{n}<\frac{(\det(X)+1)^{n}}{\det(X)^{n-1}}$. Therefore,
@@
which implies that $C(X)\cap(\mathbb{Z}^{n}\setminus\{\mathbf{0}\})\neq\emptyset$. Lemma 1 now follows from Propositions 2 and 1, and Lemma 2. ∎
@@
Proposition 5
In $n=6$, the matrix $M$ is sporadic, i.e., $C(M)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
@@
2.2 Proof of Theorem 1.1
Proof
We know that the primitive extreme points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ that acts on $\{\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}\}$. The finiteness of the index set $K$ follows from similar argument in Proposition 1. We only need to prove that the sporadic points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ on a finite set $R$.
Corollary 1 shows that for any sporadic matrix $X$, $\det(X)<\gamma_{n}$. By [21, Theorem 2.4], there exists a constant $\alpha_{n}>0$ depending only on $n$, such that for any positive definite matrix $X\in\mathcal{S}^{n}(\mathbb{Z})$, there is a unimodularly equivalent matrix $X^{\prime}$ of $X$ with diagonal entries satisfy

### v2

@@ 2 The Positive Semidefinite (PSD) Cone
The group $\mathrm{GL}(n,\mathbb{Z})$ embeds into $\mathrm{GL}(N,\mathbb{Z})$ as follows. Given a matrix $U\in\mathrm{GL}(n,\mathbb{Z})$ and any $X\in\mathcal{S}^{n}(\mathbb{Z})$, we define the action $U\cdot X:=UXU^{\mathsf{T}}$. This action is a linear map and takes integer points in $\mathbb{Z}^{N}$ to integer points, and thus can be represented by the multiplication with a matrix in $\mathrm{GL}(N,\mathbb{Z})$. It is well-known that this group $\mathrm{GL}(n,\mathbb{Z})$ is finitely generated [24]. For simplicity, we keep the notation $U\cdot X$ for the action of $U\in\mathrm{GL}(n,\mathbb{Z})$ on $X$ without explicitly specifying its representation in $\mathrm{GL}(N,\mathbb{Z})$.
@@ 2.1 Lemmas for $n\leq 5$ and $n=6$
Proof.
If there is no sporadic point in $\mathcal{S}_{+}^{n}(\mathbb{Z})$, then for every $Y\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, there exists $\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\mathbf{0}\}$ such that $Y-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0$. For $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, we do the following procedure for $X_{0}:=X$ (with index $i$ initialized to 1).
  1. 1.
    Take any $\mathbf{x}_{i}\in\mathbb{Z}^{n}\setminus\{\mathbf{0}\}$ such that $X_{i}:=X_{i-1}-\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\succeq 0$.
  2. 2.
    If $X=0$, then we have found an integer rank-1 decomposition $X=\sum_{j=1}^{i}\mathbf{x}_{j}\mathbf{x}_{j}^{\mathsf{T}}$; otherwise set the index $i\leftarrow i+1$ and go back to step 1.
To see that the procedure terminates in finitely many steps, note that the diagonal of $\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$ contains at least 1 nonzero entry because $\mathbf{x}_{i}\neq 0$. Thus the trace $\mathrm{tr}(X_{i})\leq\mathrm{tr}(X_{i-1})-1$ for any $i\geq 1$ because the entries are integers. The procedure can repeat no more than $\mathrm{tr}(X)$ times as $\mathrm{tr}(X_{i})\geq 0$.
If every $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$ has an integer rank-1 decomposition $X=\sum_{i\in K}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$, then any of $\mathbf{x}_{i},i\in K$ satisfies the requirement $X-\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\succeq 0$. ∎
For any matrix $X\in\mathcal{S}^{n}(\mathbb{R})$, we can define a convex set $K(X):=\{\mathbf{x}\in\mathbb{R}^{n}:X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0\}$. Since $X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0$ if and only if, for any $\mathbf{v}\in\mathbb{R}^{n}$, $|\mathbf{v}^{\mathsf{T}}\mathbf{x}|^{2}\leq\mathbf{v}^{\mathsf{T}}X\mathbf{v}$, we see that $K(X)$ is a compact convex set that is symmetric about the origin. This provides another equivalent formulation of the integer rank-1 decomposition.
Proposition 7.
For $X\in\mathcal{S}_{+}^{n}(\mathbb{Z})$, $X$ is sporadic if and only if $K(X)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
This provides a geometric perspective to our problem. Note that the set $K(X)$ is an ellipsoid (that may not be full-dimensional) because
@@
by the positive semidefiniteness of Schur complements, where $X^{\dagger}$ denotes the pseudoinverse of $X$. In the case where $X$ has full rank,
$K(X)=\{\mathbf{x}\in\mathbb{R}^{n}:\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}\text{ with }K(X))=V_{n}\sqrt{\det(X)}$
where $\mathrm{adj}(X)$ is the adjugate of $X$ satisfying $\mathrm{adj}(X)=\det(X)X^{-1}$ and $V_{n}:=\pi^{n/2}/\Gamma(\frac{n}{2}+1)$ is the volume of the unit $n$-ball. Note that $K(X)$ is not necessarily full-dimensional if $X$ is rank deficient. The degenerate case for rank 1 is characterized by the following proposition.
Proposition 8.
@@
Proof.
As $X\succeq 0$ and $\mathrm{rank}(X)=1$, we can assume that $X=\mathbf{a}\mathbf{a}^{\mathsf{T}}$ for some $\mathbf{a}\in\mathbb{R}^{n}$. Because $X\in\mathcal{S}^{n}(\mathbb{Z})$, we have $a_{i}a_{j}\in\mathbb{Z}$ for $i,j\in[n]$. In particular, $a_{i}^{2}\in\mathbb{Z}$. Denote $k_{i}:=a_{i}^{2}\in\mathbb{Z}_{\geq 0}$. Without loss of generality, we can assume that $k_{i}\geq 1$, i.e., $a_{i}\neq 0$, otherwise, we can just consider the submatrix corresponding to the nonzero $k_{i}$.
Suppose that there exists some $a_{i}\in\mathbb{Z}\setminus\{0\}$ and $a_{j}\in\{\pm\sqrt{k_{j}}\}\notin\mathbb{Q}$. Then $a_{i}a_{j}\notin\mathbb{Q}$, a contradiction. Therefore, we must have either $a_{i}\in\mathbb{Z}$ for all $i$ or $a_{i}\notin\mathbb{Q}$ for any $i$.
If $a_{i}\in\mathbb{Z}$ for all $i$, then the result holds with $\lambda=1$ and $x=a$.
If $a_{i}\notin\mathbb{Q}$ for all $i$, i.e., $k_{i}$ is not a square. Because $a_{i}a_{j}\in\mathbb{Z}$, we have $\sqrt{k_{i}k_{j}}\in\mathbb{Z}$, which implies that $k_{i}k_{j}=t_{ij}^{2}$ for some integer $t_{ij}$. Suppose that $p_{1},\dots,p_{s}$ are all the prime factors in the decompositions of $k_{i}$, $i\in[n]$ such that $k_{i}=\prod_{\ell=1}^{s}p_{\ell}^{\alpha_{\ell}^{i}}$, for some $\alpha_{\ell}^{i}\in\mathbb{Z}_{\geq 0}$. We have $k_{i}k_{j}=\prod_{\ell=1}^{s}p_{\ell}^{\alpha_{\ell}^{i}+\alpha_{\ell}^{j}}=t_{ij}^{2}$, which implies that $\alpha_{\ell}^{i}+\alpha_{\ell}^{j}$ is even. Therefore, for a fixed $\ell$, either $\alpha_{\ell}^{i}$ is even for all $i\in[n]$ or $\alpha_{\ell}^{i}$ is odd for all $i\in[n]$. Let $I:=\{\ell\in[s]:\alpha_{\ell}^{i}~{}\text{is odd}\}$ and $\lambda=\prod_{\ell\in I}p_{\ell}$. We have $k_{i}/\lambda$ is a square, thus $X=\lambda\mathbf{x}\mathbf{x}^{\mathsf{T}}$ for $x=a/\sqrt{\lambda}$, where $x_{i}=\sqrt{k_{i}/\lambda}\in\mathbb{Z}$. ∎
From Proposition 8, we can directly prove the case for $n=2$ using Minkowski’s First Theorem (for example, see [6]): if $\mathrm{vol}(K(X))>4$, then there is a nonzero integral point in $K(X)$.
Proposition 9.
@@
Proof.
For $n=2$. Let $X=\begin{bmatrix}a_{11}&a_{12}\\ a_{12}&a_{22}\end{bmatrix}\succeq 0$, where $a_{11},a_{12},a_{22}\in\mathbb{Z}$, which implies that $a_{11}\geq 0,a_{22}\geq 0,a_{11}a_{22}-a_{12}^{2}\geq 0$. By Proposition 8, we only need to consider the case when $X\succ 0$, i.e., $\det(X)=a_{11}a_{22}-a_{12}^{2}\geq 1,a_{11}\geq 1,a_{22}\geq 1$. Then,
$K(X)$ :=$:=\{\mathbf{x}\in\mathbb{R}^{2}:X-\mathbf{x}\mathbf{x}^{\mathsf{T}}\succeq 0\}$
=$=\{\mathbf{x}\in\mathbb{R}^{2}:a_{11}-x_{1}^{2}\geq 0,a_{22}-x_{2}^{2}\geq 0,(a_{11}-x_{1}^{2})(a_{22}-x_{2}^{2})-(a_{12}-x_{1}x_{2})^{2}\geq 0\}.$
We claim that $K(X)=\{\mathbf{x}\in\mathbb{R}^{2}:(a_{11}-x_{1}^{2})(a_{22}-x_{2}^{2})-(a_{12}-x_{1}x_{2})^{2}\geq 0\}.$ To see this, note that
$0\leq(a_{11}-x_{1}^{2})(a_{22}-x_{2}^{2})-(a_{12}-x_{1}x_{2})^{2}$ $=(a_{11}a_{22}-a_{12}^{2})-a_{22}x_{1}^{2}+2a_{12}x_{1}x_{2}-a_{11}x_{2}^{2}$
$=\frac{a_{11}a_{22}-a_{12}^{2}}{a_{11}}(a_{11}-x_{1}^{2})-a_{11}(x_{2}-\frac{a_{12}}{a_{11}}x_{1})^{2}$
which together with $a_{11},a_{22},a_{11}a_{22}-a_{12}^{2}>0$ implies that $a_{11}-x_{1}^{2}\geq 0$. Similarly,
$0\leq(a_{11}-x_{1}^{2})(a_{22}-x_{2}^{2})-(a_{12}-x_{1}x_{2})^{2}=\frac{a_{11}a_{22}-a_{12}^{2}}{a_{22}}(a_{22}-x_{2}^{2})-a_{22}(x_{1}-\frac{a_{12}}{a_{22}}x_{2})^{2}$
implies that $a_{22}-x_{2}^{2}\geq 0$, which shows the claim.
Since $X$ has full rank, $K(X)=\{\mathbf{x}\in\mathbb{R}^{2}:\det(X)-a_{22}x_{1}^{2}+2a_{12}x_{1}x_{2}-a_{11}x_{2}^{2}\geq 0\}$ is a centrally symmetric ellipsoid with area $\pi\sqrt{\det(X)}$.
If $\det(X)\geq 2$, then $\mathrm{vol}(K(X))\geq\sqrt{2}\pi>4$, we know that $K(X)\cap\mathbb{Z}^{2}\neq\emptyset$ by Minkowski’s First Theorem.
If $\det(X)=1$, then $K(X)=\{x\in\mathbb{R}^{n}:a_{22}x_{1}^{2}-2a_{12}x_{1}x_{2}+a_{11}x_{2}^{2}\leq 1\}$. Because $a_{22},a_{12},a_{11}\in\mathbb{Z}$, we have $K(X)\cap\mathbb{Z}^{2}=\tilde{K}(X)\cap\mathbb{Z}^{2}$, where $\tilde{K}(X)=\sqrt{2-\epsilon}\cdot K(X)=\{x\in\mathbb{R}^{n}:a_{22}x_{1}^{2}-2a_{12}x_{1}x_{2}+a_{11}x_{2}^{2}\leq 2-\epsilon\}$ for any $10<\epsilon<$ 1. Therefore, $\mathrm{vol}(\tilde{K}(X))=(2-\epsilon)\cdot\pi>4$ when we take any $\epsilon<2-\frac{4}{\pi}$. Therefore, $\tilde{K}(X)\cap\mathbb{Z}^{2}$ is nonempty by Minkowski’s First Theorem so neither is $K(X)\cap\mathbb{Z}^{2}$. ∎
In the degenerate case, we can reduce the problem to one involving full-rank matrices of some lower dimension.
Lemma 10.
@@
Proof.
If $\mathrm{rank}(X)<n$, then there exists a primitive vector $\mathbf{z}\in\mathbb{Z}^{n}$ in the subspace $N:=\{\mathbf{y}\in\mathbb{R}^{n}:~{}X\mathbf{y}=\bf 0\}$. The sublattice $\Lambda:=\mathbb{Z}^{n}\cap N$ is primitive and thus a basis of $\Lambda$ containing $\mathbf{z}$ can be extended to a basis of $\mathbb{Z}^{n}$, which we denote as $U=[\mathbf{z},\mathbf{u}_{2},\dots,\mathbf{u}_{n}]$ [25, Chapter 2, Lemma 4]. Because $U$ is a basis of $\mathbb{Z}^{n}$, we know that $|\det(U)|=1$, i.e., $U$ is unimodular. Then
$U^{\mathsf{T}}XU=\begin{bmatrix}0&0\\ 0&\hat{X}\end{bmatrix}.$
Iterating this process until $\hat{X}$ is positive definite, i.e., $\mathrm{rank}(\hat{X})=r$. ∎
Notice that if $X_{1},X_{2}$ are unimodularly equivalent, then $K(X_{1})\cap\mathbb{Z}^{n}\neq\{\mathbf{0}\}$ if and only if $K(X_{2})\cap\mathbb{Z}^{n}\neq\{\mathbf{0}\}$. Thus our problem expects an answer under the unimodular equivalence of integer matrices in $\mathcal{S}_{+}^{n}(\mathbb{Z})$.
The scaling of $K(X)$ into $\tilde{K}(X)$ (while preserving the integer points) in the proof for Proposition 9 results in
$\tilde{K}(X))<V_{n}\sqrt{\det(X)}\cdot\biggl{(}\frac{\det(X)+1}{\det(X)}\biggr{)}^{n/2}$
where $V_{n}$ is a constant dependent only on the dimension $n$, and the right-hand side can be approached arbitrarily. When $n=3$, $V_{n}\approx 4.189$, the right-hand side becomes $2^{3/2}\approx 2.828$, $\sqrt{2}\cdot(3/2)^{3/2}\approx 2.598$, $\sqrt{3}\cdot(4/3)^{3/2}\approx 2.667$ for $\det(X)=1,2,$ and $3$, respectively, and greater than 2 for $\det(X)\geq 4$. Thus $\tilde{K}(X))>8$ so $\tilde{K}(X)\cap\mathbb{Z}^{3}\neq\{0\}$ by Minkowski’s First Theorem.
@@
$\gamma_{n}:=\left(\max_{A\succ 0}\frac{\lambda_{1}(A)}{(\det(A))^{\frac{1}{n}}}\right)^{n},\text{ where }\lambda_{1}(A)=\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}A\mathbf{x}).$
Remark 1.
@@
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $24$
$\gamma_{n}$ $\frac{4}{3}$ $2$ $2$ $8$ $\frac{64}{3}$ $64$ $256$ $4^{24}$
Remark 2.
From a volume argument by Minkowski’s First Theorem, we have
$\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}A\mathbf{x})\leq\frac{4}{\pi}\Gamma(1+\frac{n}{2})^{\frac{2}{n}}\det(A)^{\frac{1}{n}}\sim\frac{2n}{\pi e}\det(A)^{\frac{1}{n}},$
which is better than the bound given by $\gamma_{n}\leq(\frac{4}{3})^{\frac{n(n-1)}{2}}$ when $n$ is large. But this estimation on $\gamma_{n}$ is not enough to prove Lemma 5 for the dimension $n=4,5$.
Proof of Lemma 5.
The case $n=1$ follows from Proposition 8. We will show that $K(X)\cap\mathbb{Z}^{n}\neq\{\mathbf{0}\}$ for $2\leq n\leq 5$, where $K(X)=\{\mathbf{x}\in\mathbb{R}^{n}:\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq\det(X)\}$. By the definition of the Hermite constant, we have
@@
For $n=2,3,4,5$, we have $\frac{n^{n}}{(n-1)^{n-1}}>\gamma_{n}$ as $\frac{2^{2}}{1^{1}}=4$, $\frac{3^{3}}{2^{2}}\approx 6.75$, $\frac{4^{4}}{3^{3}}\approx 9.48$, $\frac{5^{5}}{4^{4}}\approx 12.21$, and $\frac{6^{6}}{5^{5}}\approx 14.93$. By taking the derivative with respect to $\det(X)$ for finding extremum value, we know that $\frac{(\det(X)+1)^{n}}{(\det(X))^{n-1}}\geq\frac{n^{n}}{(n-1)^{n-1}}$. Thus, $\gamma_{n}<\frac{(\det(X)+1)^{n}}{\det(X)^{n-1}}$. Therefore,
@@
which implies that $K(X)\cap(\mathbb{Z}^{n}\setminus\{\mathbf{0}\})\neq\emptyset$. Lemma 5 now follows from Propositions 7 and 6, and Lemma 10. ∎
@@
Proposition 12.
In $n=6$, the matrix $M$ is sporadic, i.e., $K(M)\cap\mathbb{Z}^{n}=\{\mathbf{0}\}$.
@@
Proof.
We verify that $\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})>\det(X)=3$.
$\mathrm{adj}(X)=(\det(X))X^{-1}=\begin{bmatrix}4&3&1&-2&-2&-2\\ 3&6&3&-3&-3&-3\\ 1&3&4&-2&-2&-2\\ -2&-3&-2&4&1&1\\ -2&-3&-2&1&4&1\\ -2&-3&-2&1&1&4\end{bmatrix}$
Note that $\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}=[(x_{1}+2x_{2}+x_{3}-x_{4}-x_{5}-x_{6})^{2}+(x_{1}+x_{2}+x_{3}-x_{4}-x_{5}-x_{6})^{2}+x_{2}^{2}]+[(x_{1}-x_{3})^{2}+x_{1}^{2}+x_{3}^{2}]+[(x_{4}-x_{5})^{2}+(x_{4}-x_{6})^{2}+(x_{5}-x_{6})^{2}]$. Let $A_{2}:=(x_{1}+2x_{2}+x_{3}-x_{4}-x_{5}-x_{6})^{2}+(x_{1}+x_{2}+x_{3}-x_{4}-x_{5}-x_{6})^{2}+x_{2}^{2}$, $A_{13}:=(x_{1}-x_{3})^{2}+x_{1}^{2}+x_{3}^{2}$ and $A_{456}:=(x_{4}-x_{5})^{2}+(x_{4}-x_{6})^{2}+(x_{5}-x_{6})^{2}$. Then $\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}=A_{2}+A_{13}+A_{456}$.
Suppose, there exists $x\in\mathbb{Z}^{6}$ such that $\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}\leq 3$. We are going to show that $x=0$. Notice that $A_{2},A_{13},A_{456}$ are even, which implies that $A_{2}+A_{13}+A_{456}\leq 2$. Then at most one of $A_{2},A_{13},A_{456}$ is nonzero.
We consider the following three cases:
  1. 1.
    if $A_{13}=0,A_{456}=0$, then $x_{1}=x_{3}=0,x_{4}=x_{5}=x_{6}=0$. Because $A_{2}=6x_{2}^{2}\leq 2$, we have $x_{2}=0$.
  2. 2.
    if $A_{2}=0,A_{456}=0$, then $x_{4}=x_{5}=x_{6}=0$, $x_{2}=0$, $x_{1}+x_{3}=0$. Because $A_{13}=6x_{1}^{2}\leq 2$, we have $x_{1}=0$.
  3. 3.
    if $A_{2}=0,A_{13}=0$, then $x_{1}=x_{3}=0$, $x_{2}=0$, $x_{4}+x_{5}+x_{6}=0$. Because $A_{456}=(x_{4}-x_{5})^{2}+(2x_{4}+x_{5})^{2}+(x_{4}+2x_{5})^{2}=6(x_{4}^{2}+x_{5}^{2}+x_{4}x_{5})\leq 2$, we have $x_{4}=x_{5}=x_{6}=0$.
Therefore, $\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})>3$. For $\mathbf{x}=\mathbf{e}_{1}$, $\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x}=4$, i.e., $\min_{\mathbf{x}\in\mathbb{Z}^{n}\setminus\{\bf 0\}}(\mathbf{x}^{\mathsf{T}}\mathrm{adj}(X)\mathbf{x})=4$. ∎
@@
2.2 Proof of Theorem 1
Proof.
We know that the primitive extreme points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ that acts on $\{\mathbf{e}_{1}\mathbf{e}_{1}^{\mathsf{T}}\}$. Thus we only need to prove that the sporadic points are generated from the group $\mathrm{GL}(n,\mathbb{Z})$ on a finite set $R$.
Corollary 11 shows that for any sporadic matrix $X$, $\det(X)<\gamma_{n}$. By [26, Theorem 2.4], there exists a constant $\alpha_{n}>0$ depending only on $n$, such that for any positive definite matrix $X\in\mathcal{S}^{n}(\mathbb{Z})$, there is a unimodularly equivalent matrix $X^{\prime}$ of $X$ with diagonal entries satisfy